A body covers 12m in 2nd second and 20m in 4th second. Find what distance the body will cover in 4 seconds after the 5th second .

Grade:12th pass

1 Answers

Eshan
5 years ago
Dear student,

Distance travelled in nth second iss_n=u+\dfrac{a}{2}(2n-1)
Hence12=u+\dfrac{3a}{2}

and20=u+\dfrac{7a}{2}

\implies a=4m/s^2andu=6m/s

Hence distance convered in 4 seconds after 5thsecond=s(t_2=9)-s(t_1=5)=u(t_2-t_1)+\dfrac{1}{2}a(t_2^2-t_1^2)=136m